One alloy of iron consists 49% iron & another 28% irob.How many tons should be used to make 120 tons of 42% iron alloy?
Answer by lwsshak3(11628)

One alloy of iron consists 49% iron & another 28% irob.How many tons should be used to make 120 tons of 42% iron alloy?
***
let x=amt of 49% iron to use
120-x=amt of 28% iron to use
..
49%x+28%(120-x)=42%*120
.49x+33.6-.28x=50.4
.21x=16.8
x=80
120-x=40
amt of 49% iron to use=80 tons
amt of 28% iron to use=40 tons